### 27 Nov 2023 v1

# Have we built machines that think like people?

Abstract
A chief goal of artificial intelligence is to build machines that think like people. Yet it has been argued that deep neural network architectures fail to accomplish this. Researchers have asserted these models’ limitations in the domains of causal reasoning, intuitive physics, and intuitive psychology. Yet recent advancements, namely the rise of large language models, particularly those designed for visual processing, have rekindled interest in the potential to emulate human-like cognitive abilities. This paper evaluates the current state of vision-based large language models in the domains of intuitive physics, causal reasoning, and intuitive psychology. Through a series of controlled experiments, we investigate the extent to which these modern models grasp complex physical interactions, causal relationships, and intuitive understanding of others’ preferences. Our findings reveal that, while these models demonstrate a notable proficiency in processing and interpreting visual data, they still fall short of human capabilities in these areas. The models exhibit a rudimentary understanding of physical laws and causal relationships, but their performance is hindered by a lack of deeper insights—a key aspect of human cognition. Furthermore, in tasks requiring an intuitive theory of mind, the models fail altogether. Our results emphasize the need for integrating more robust mechanisms for understanding causality, physical dynamics, and social cognition into modern-day, vision-based language models, and point out the importance of cognitively-inspired benchmarks.
Introduction
Figure 1: Overview of domains tasks, approach, and models. A: Example images for the different experiments. Each experiment was taken from one of three cognitive domains: intuitive physics, causal reasoning, and intuitive psychology. B: General approach. For every query, an image was submitted to the model, and different questions were asked about the image, i.e. we performed visual question answering.C: Used multi-modal large language models and their size.
People are quick to anthropomorphize, attributing human characteristics to non-human agents [1]. This tendency, far from being a modern quirk, is deeply rooted in our collective psyche, as evidenced through our literary traditions. Consider E.T.A. Hoffmann’s “Der Sandmann” [2], where Nathaniel’s infatuation with Olimpia, a mere mechanical automaton with a limited repertoire of expressions – her most notable being a plaintive ’Ach, Ach!’ – serves as a poignant example. With further technological advances, this phenomenon, once confined to the realm of literature, has now manifested in reality. Already simple chatbots like ELIZA [3], which used simple pattern matching and substitution methodologies, gave users the illusion of human-like understanding and engagement [4]. The tendency to anthropomorphize has only intensified with the advent of Large Language Models (LLMs) [5]. These advanced AI systems represent a big leap from ELIZA both in complexity and capability. Unlike ELIZA’s simplistic pattern matching, LLMs apply deep learning techniques to generate text [6], learning from vast datasets to produce responses that can be startlingly human-like [7]. Astonishingly, these models cannot only generate text. When scaled up to bigger training data and architectures, other, so-called “emergent abilities” appear [8, 9]. The current models can, for example, pass the bar exam [10], write poems [11], compose music [12], and assist in programming and data analysis tasks [13]. As a result, the line between human and machine capabilities is increasingly blurred [14, 15]. People not only interact with these systems as if they were humans [16], but they also start to rely on them for complex decision-making [17], artistic creation [18], and personal interactions [19]. It is, therefore, natural to ask: Have we built machines that think like people? Or are we, just like Nathaniel, projecting human qualities into rudimentary entities that are fundamentally different from us?
Judging whether or not artificial agents can mimic human thought is at the core of cognitive science [20, 21]. Therein, researchers have long debated the capabilities of artificially intelligent agents [22, 23, 24]. In a seminal paper, Lake et al. [25] proposed core domains to consider when making such judgments. Published during the height of the deep learning revolution [26], the authors focused on domains that were easy for people but difficult for deep learning models: intuitive physics, causal reasoning, and intuitive psychology.
Research on intuitive physics has studied how people perceive and interpret physical phenomena [27, 28, 29]. Past work on this topic has emphasized that humans possess an innate ability to predict and understand the physical properties of objects and their interactions [30], even from a young age [31], a notion sometimes summarized as a “physics engine” in people’s heads [32]. This understanding includes concepts such as gravity [33], inertia [34], and momentum [35]. Some of the most canonical tasks in this domain involve testing people’s judgments about the stability of block towers [36, 37]. These tasks have made their way into machine learning benchmarks [38, 39], where they are used to test the intuitive physical understanding of neural networks (see [40] for an overview of previous work on building models with human-like physical knowledge).
Research on causal reasoning has studied how individuals infer and think about cause-effect relationships [41, 42, 43]. Past work on this topic has proposed that humans possess an intuitive capacity to infer, understand, and predict causal relationships in their environment [44, 45, 46, 47], oftentimes described using Bayesian models of causal learning [48, 49]. This cognitive ability encompasses recognizing patterns [50, 51], inferring causes from interventions [52, 53], and predicting future events based on hypothetical events [54]. Canonical tasks in this domain often involve assessing individuals’ ability to infer causal relationships, for example, when judging the responsibility of one object causing other objects’ movement [55, 56]. Causal reasoning remains a challenge, even for current machine learning approaches [57, 58].
Research on intuitive psychology has explored how individuals infer, understand, and interpret social phenomena and mental states of other agents [59, 60]. Past work on this topic has emphasized the concept that humans possess an inherent ability to infer and reason about the mental states [61, 62], intentions, and emotions of others, often referred to as a “theory of mind” [63, 64]. This ability has been modeled as a Bayesian inference problem [65, 66, 67]. Canonical tasks in this domain often involve assessing individuals’ capacity to predict actions based on understanding others’ perspectives or intentions, such as determining agents’ utility functions based on their actions in a given environment [68, 69]. It is the subject of ongoing debates if modern algorithms show any form of intuitive psychology [70, 71, 72].
Lake et al. argued that some of these abilities act as “start-up software,” because they constitute cognitive capabilities present early in development. Moreover, they proposed that these so-called “intuitive theories” [73, 74] need to be expressed explicitly using the calculus of Bayesian inference [75], as opposed to being learned from scratch, for example, via gradient descent. However, with the increase in abilities of current neural networks, in particular LLMs, we pondered: Can LLMs, in particular vision LLMs, sufficiently solve problems from these core domains?
To address this question, we took canonical tasks from the domains of intuitive physics, causal reasoning, and intuitive psychology that could be studied by providing images and language-based questions. We submitted them to some of the currently most advanced LLMs. Our results showed that these models can, at least partially, solve these tasks. In particular, the largest currently available model, OpenAI’s Generative Pre-trained Transformer (GPT-4) managed to perform robustly above chance in two of the three domains. Yet crucial differences emerged. First, none of the models matched human-level performance in any of the domains. Secondly, none of the models fully captured human behavior, leaving room for domain-specific models of cognition such as the Bayesian models originally proposed for the tasks.
Results
We tested four different models on three core components for human-like intelligence as outlined by Lake et al. ([25]; see Fig. 1A). The models we used are vision large language models, which are multimodal models that integrate image processing capabilities into large language models ([76, 77]; see Fig. 1B). These models allow users to perform visual question answering [78, 79]: users can upload an image and ask questions about it, which the model interprets and responds to accordingly.
The first model was GPT-4 with Vision (GPT-4V), developed by OpenAI [80]. This multimodal model extends the abilities of GPT-4 to analyze and interpret images, although the details of how the model accomplished this have not been made public. The second model is ADEPT’s Fuyu-8B [81], which is a decoder-only multimodal model based on the classic transformer architecture. Fuyu-8B stands out due to its simpler architecture and training procedure compared to other multimodal models. The third model is LLaMA-Adapter V2 [82], or short Adapter, which offers an alternative approach to enhance the capabilities of vision-language models by increasing learnable parameters, introducing an early fusion strategy for better visual integration, and employing a joint training approach for image-text and instruction-following data. The final model is Otter [83], which is based on OpenFlamingo [84], an open-sourced version of DeepMind’s Flamingo [85], and specifically designed for multimodal in-context instruction following.
For testing the three core components, we used three tasks taken from the cognitive science literature that could be studied in vision LLMs via visual question answering. For intuitive physics, we asked models to judge the stability of different block towers, using stimuli originally proposed by Lerer and colleagues [86] and based on previous research by Battaglia et al. [36]. For causal reasoning, we again used block towers and asked how many blocks would fall if certain blocks were removed or to judge the responsibility of certain blocks for the stability of the tower, adopting a design that Zhou and colleagues had previously used in human subjects [87]. Finally, for intuitive psychology, we used a task where the models saw a picture of an agent’s path on a grid and then had to make inferences about the costs and rewards associated with the environment, taken from Jara-Ettinger and colleagues [88].
For every task, we queried the visual reasoning abilities of the LLMs with tasks of increasing complexity (see Fig. 1C). First, we asked about simple features of the shown images such as the background color or the number of objects shown. Afterward, we submitted questions taken from the cognitive science experiments. We report results based on comparisons to the ground truth as well as the different models’ matches to human data.
Figure 2: Results for four different large language models for tasks of increasing complexity given images of real block towers from [86]. We first ask for the background color in the image (A), then the number of blocks in the image (B), and finally a binary stability rating for the block towers (C). The last plot shows the match to human stability ratings as determined by the Pearson correlation (D). Errors bars in plots A - C are given by the standard deviation of a binomial distribution. The error bars in plot D are given by the 95% confidence interval for the correlation coefficient.
Intuitive physics
To test the intuitive physics capabilities of the different LLMs, we used photographs depicting wooden block towers from Lerer et al. [86] (see Fig. 5 in the Appendix for an example). These images mirror stimuli that developmental psychologists use to study the development of intuitive physics in infants. We used these images to test the models in increasingly complex tasks, starting with determining the background color of a given image (1), counting the number of colored blocks in the image (2), and giving a binary stability judgment of the depicted block tower (3).
All four models were able to correctly perform the first and easiest task: they all achieved almost perfect accuracy in determining the background color of the images (see Fig. 2A). In the second task, the performance of most models deteriorated, with only GPT-4V correctly determining the number of blocks in all images (see Fig. 2B). It is important to note that the first two tasks are fairly trivial for humans and we would expect human performance to be at 100% (the background color is always white and images featured either 2, 3, or 4 blocks).
[…]
Causal reasoning
To test the causal reasoning capabilities, we used synthetic images again depicting block towers from Zhou et al. [89, 87] (see Fig. 6 in the Appendix for an example). Here, the images showed static scenes of block towers that were stable but might collapse if one of the blocks was removed.
Figure 3: Results for analyses using images from causal reasoning experiments taken from [89]. We first ask for the number of blocks in the image, then the number of blocks that would fall if a specific block is removed, and finally a rating between 0 and 100 for how responsible a specific block is for the stability of the tower. For the responsibility ratings, all LLMs except for GPT-4V give constant ratings: Fuyu always responds with 100, while Otter and LLaMA-Adapter V2 always respond with 50. Error bars in plots A - C are given by the standard error of the mean, while the error bar plot D is given by the 95% confidence interval of the correlation coefficient.
Again, we started by asking the models to count the blocks in the image (1), we continued by querying the models for the number of blocks that would fall if a specific block was removed from the scene (2), and finally, we asked the models to rate the responsibility of a specific block for the stability of the other blocks (3). For the second task, we established a baseline performance represented by a horizontal line in Fig. 3A-C, which corresponds to a random agent. This random agent used a simple strategy: it gave the mean between 0 and the number of blocks in each image as its’ prediction, essentially behaving like a uniform distribution over the possible number of blocks that could fall.
The images in this task displayed a larger number of blocks (ranging from 6 to 19), which made the basic counting task significantly more challenging than in the previous section. Models’ responses approximated the ground truth, albeit rarely matching it exactly. Therefore, we report the mean absolute distance to the ground truth instead of the percentage of correct answers (see Fig. 3A). The models’ performance highlighted the challenging nature of this task, with the best performing model (GPT4-V) still being on average  2 blocks off.
[…]
Intuitive psychology
To test the intuitive psychology of the different LLMs we used synthetic images depicting an astronaut on a colored background from Jara-Ettinger et al. [88] (see Fig. 5 in the Appendix for an example). The experiment consisted of three parts which differed slightly in their layout and components. Depending on the experiment, the images featured either one or two different terrains (indicated by different background colors and textures) as well as either one or two different care packages. The astronaut was shown with a path that led from a starting point to a base; the astronaut could collect care packages along the way. Depending on which terrain the astronaut crossed or which care package they chose to pick up or not, it was possible to infer the costs associated with the different terrains and the rewards associated with the different care packages.
Figure 4: Results on tasks for intuitive psychology taken from [88]. Again, we first ask for the background color (A). Models are then asked to make inferences about the costs and rewards in an environment depending on the path an agent has taken (B and C). Correlation values for Fuyu and LLaMA-Adapter V2 are missing as they always responded with constant ratings for either cost or reward questions. Error bars in plot A are given by the standard deviation of a binomial distribution, while the error bars in plots B and C are given by the 95% confidence interval for the correlation coefficient.
[…]
Discussion
We started by asking whether, with the rise of modern large language models, researchers have created machines that – at least to some degree – think like people. To address this question, we took four recent multi-modal large language models and probed their abilities in three core cognitive domains: intuitive physics, causal reasoning, and intuitive psychology.
In intuitive physics, the models managed to solve some of the given tasks and showed a medium match with human data. Similarly, in a causal reasoning task, some models, in particular, GPT-4V, performed well and showed a medium match with human data. Finally, in an intuitive psychology task, none of the models performed well and none of them showed a reasonable match with human data. Thus, an appropriate answer to the question motivating our work would be “No.”, or – perhaps more optimistically – “Not quite.”
Limitations and Future Work
Although we have tried our best to give all models a fair chance and set up the experiments in a clean and replicable fashion, some shortcomings remain that should be addressed in future work.
First of all, we have only tested a handful of multi-modal models on just three cognitive domains. While we believe that the used models and tasks provide good insights into the state-of-the-science of LLMs’ cognitive abilities, future studies should look at more domains and different models to further tease apart when and why LLMs can mimic human reasoning. For example, it would be interesting to see if scale is the only important feature influencing model performance [90, 91]. Currently, our evidence suggests that even smaller models, for example Fuyu with its 8 billion parameters, can sometimes perform as well as GPT-4V in some tasks.
Another shortcoming of the current work is the simplicity of the used stimuli. While the block towers used in our first study were deliberately designed to be more realistic [86] than commonly used psychological stimuli [36], this was not true for the experiments in the other two domains. For the intuitive psychology experiments, in particular, we would expect the models to perform better if the stimuli contained more realistic images of people, which has been shown to work better in previous studies [92]. Interestingly, using more realistic stimuli can also change people’s causal judgments [93]; how realistic stimuli used in cognitive experiments should be, remains an open question [94].
On a related point, we only used static images in our current experiments, which severely limits the breadth and level of detail of the questions we could ask. For example, some of the most canonical tasks investigating people’s causal reasoning abilities involve videos of colliding billiard balls [55]. As future large language models will likely be able to answer questions about videos [95], these tasks represent the next frontier of cognitively-inspired benchmarks.
For the comparisons to human data, we currently used the participant data collected in the original studies and assessed the correspondence between models and this data via correlation coefficients. Future work could expand on this approach by collecting new data from human subjects choosing which of the model’s judgments they prefer. This could lead to a more detailed comparison, similar to what has been proposed to discriminate among deep learning models for human vision [96] and language [97].
A crucial weakness of most studies using large language models is that they can be sensitive to specific prompts [98, 99, 100]. While we have attempted to use prompts that elicited good behavior, thereby giving LLMs a chance to perform well, future work could try to further optimize these prompts using available methods [101, 102, 103], while also assessing how the models respond to paraphrased versions of the same tasks. While it could be possible to further engineer the used prompts, as well as try out several other ways of phrasing the same prompt, we believe that our current approach was sufficient to showcase these models’ abilities.
Finally, we applied all models out of the box and without additional fine-tuning. Future studies could attempt to fine-tune multi-modal LLMs to better align with cognitive data [104] and assess if this improves their reasoning abilities more generally.
Related work
We are not the first to assess LLMs’ reasoning abilities [105, 106, 107]. Previous studies have focused, among others, on testing LLMs’ cognitive abilities in model-based planning [108], analogical reasoning tests [109], exploration tasks [110], systematic reasoning tests [111, 112], psycholinguistic completion studies [113], and affordance understanding problems [114]. In this sense, our contribution can be seen as a part of a larger movement in which researchers use methods from the behavioral sciences to understand black box machine learning models [115, 116, 117]. However, most of the previous studies did not investigate multi-modal LLMs but rather remained in the pure language domain. Although there are recent attempts to investigate vision LLMs cognitive features, including their reaction to visual illusions [118] as well as how they solve simple intelligence tasks [119], we are the first to investigate the proposed core components of cognition in these models.
Previous work has also looked at how LLMs solve cognitive tasks taken from the same domains that we have looked at. In intuitive physics, Zečević et al. [120] found that LLMs performed poorly in a task using language descriptions of physical scenarios. Zhang and colleagues [121] extracted programs from text produced by large language models to improve their physical reasoning abilities. Finally, Jassim and colleagues [122] proposed a novel benchmark for evaluating multimodal LLMs’ understanding of situated physics. In causal reasoning, Binz and Schulz [123] showed that GPT-3 failed at simple causal reasoning experiments, while Kosoy et al. [124] showed that LLMs cannot learn human-like causal over-hypotheses. In research on intuitive psychology, Kosinsky argued that theory of mind might have emerged in LLMs [71] which has been criticized other researchers [72]. Akata et al. showed that GTP-4 plays repeated games very selfishly and could not pick up on simple conventions such as alternating between options [19]. Finally, Gandhi and colleagues [125] proposed a framework for procedurally generating Theory of Mind evaluations and found that GPT4’s abilities mirror human inference patterns, though less reliable, while all other LLMs struggled.
Many of the past studies on LLMs have fallen risk of appearing in new models’ training set. Recent work has recognized this issue and, in turn, evaluated language models on many problem variations to minimize training set effects [126]. Our work differs from these approaches as current models could not have just memorized solutions to the given problems because these problems require deep reasoning and are rarely ever published with the ground truth attached.
Conclusion
A major plot twist in E.T.A Hoffmann’s “Der Sandmann” is that Nathaniel might have been a machine himself, which explains why he fell for Olimpia. This metaphor also relates to our anthropomorphization of LLMs: since LLMs are trained on human-generated data, their behaviors will always reflect our behaviors and biases. However, this reflection is sharpening, and modern neural network architectures have become more human-like. One of the other domains emphasized by Lake et al. was the ability to reason compositionally. Recent attempts have shown that neural networks can perform compositional reasoning if trained appropriately [127]. Similarly, our current work has shown that multimodal LLMs have come a long way, showing some correspondence to human behavior and often performing above chance. Moreover, machine learning researchers have put forward various ideas about how to close the remaining gap between humans and machines [128], including self-supervised learning [129], translating from natural into probabilistic languages [130], or grounding LLMs in realistic environments [131]. This continuous evolution in models’ capabilities necessitates a reevaluation of the metaphors and tools we use to understand them. We believe that cognitive science can offer tools, theories, and benchmarks to evaluate how close we have come to building machines that think like people.
Methods
Code
The open-source models were installed per the instructions on their related Github or Huggingface repositories and evaluated on a Slurm-based cluster with a single A100. For the results reported as GPT-4V, we used the public ChatGPT interface and the OpenAI API, specifically the November 2023 release of gpt4-vision-preview model which is available via the completions endpoint. Code for replicating our results is available on GitHub (github.com/lsbuschoff/multimodal). All models were evaluated in Python using PyTorch [132]. Additional analyses were carried out using NumPy [133], Pandas [134], and SciPy [135]. Matplotlib [136] and Seaborn [137] were used for plotting.
Models
GPT4-V
We initially queried GPT-4V through the ChatGPT interface, since the OpenAI API was not publicly available at the outset of this project. The Intuitive Psychology task responses were collected using the gpt4-vision-preview model variant after its November 2023 release in the API. We set the maximum number of generated tokens for a given prompt to 1 to get single numerical responses. All other parameters were set to their default values. Note that this model does not currently feature an option for manually setting the temperature, and the provided documentation does not specify what the default temperature is.
FUYU
Fuyu is an 8B parameter multi-modal text and image decoder-only transformer. We used the Huggingface implementation with standard settings and without further finetuning (available here). The maximum number of generated tokens was set to 8 and responses were parsed by hand.
ADAPTER
We here use the multi-modal version of LLaMA-Adapter V2, which adds adapters into LLaMA’s transformer in order to turn it into an instruction-following model. We used the GitHub implementation with standard settings and again without further finetuning (available here). The maximum number of generated tokens was left at 512 and responses were parsed by hand.
OTTER
Otter is a multi-modal LLM that supports in-context instruction tuning and it is based on the OpenFlamingo model. We used the Huggingface implementation (available here), again with standard settings and without finetuning. The maximum number of generated tokens was left at 512 and responses were parsed by hand.
Datasets
Intuitive physics
We tested the intuitive physical understanding of the models using images from Lerer et al. [86]. The photos depict a block tower consisting of colored wooden blocks in front of a white fabric (see Fig. 5 for an example). The images are of size 224 x 244. In the data set, there are a total of 516 images of block towers. We tested the models on 100 randomly drawn images. We first tested the models on their high-level visual understanding of the scenes: we tasked them with determining the background color and the number of blocks in the image. In order to test their physical understanding, we tested them on the same task as the original study: we asked them to give a binary rating on the stability of the depicted block towers. For the first two tasks, we calculated the percentage of correct answers for each of the models. For the third task, we used the Pearson correlation to determine the degree of similarity between the models and human subjects.
Causal reasoning
For the causality part we used images from Zhou et al. [87]. The images show artificial block stacks of red and gray blocks on a black table (see Fig. 6 for an example). The data set consists of 42 images on which we tested all models. We again first tested the models on their high-level visual understanding of the scene and therefore tasked them with determining the number of blocks in the scene. The ground truth number of blocks in the scenes ranged from 6 to 19. Since this task is rather challenging due to the increased number of blocks, we do not report the percentage correct as for the intuitive physics data set but the mean over the absolute distance between model predictions and the ground truth for each image (see Fig. 3A).
To test the causal reasoning of the models we adopted the tasks performed in the original study [89]. We asked models to infer how many red blocks would fall if the gray block was removed. We again report the absolute distance between model predictions and the ground truth for each image (see Fig. 3B). We calculate a random baseline which uses the mean between 0 and the number of blocks for each specific image as the prediction. We also ask the models for a rating between 0 and 100 for how responsible the gray block is for the stability of the tower. For both, the number of blocks that would fall if the gray block was removed, and its’ responsibility for the stability of the tower, we calculate the mean Pearson correlation to human subjects from the original study (see Fig. 3C).
Intuitive psychology
To test the intuitive psychology of the different LLMs, we used stimuli from Jara-Ettinger et al. [88]. This part consisted of three different experiments each consisting of 16, 17, and 14 images showing a 2D depiction of an astronaut and care packages in different terrains (see Fig. 5 for an example). In order to check their high-level understanding of the images, we again asked the models to determine the background color of the images. Since this background color is not uniform, we counted both “Pink” and “Purple” as correct answers. We report the percentage of correct answers for the background color in Figure 2A.
In accordance with the original study, analyses for the intuitive psychological capabilities of the models are split into cost questions (passing through a terrain is associated with a cost for the agent) and reward questions (collecting a care package yields some sort of reward for the agent). We pooled cost and reward questions over all three experiments and reported the mean Pearson correlation with both humans and a heuristic outlined in [88] (see Figs. 2B and 2C). This heuristic calculates the costs and rewards associated with the environment from the amount of time an agent spends in each terrain and which care package it collects.
Acknowledgements
This work was supported by the Max Planck Society, the Volkswagen Foundation, the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A, and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2064/1 – 390727645.
Author contributions statement
All authors conceived the experiments. L.B.S. and E.A. conducted the experiments. All authors analysed the results. All authors wrote the manuscript.
Appendix A Example trials
A.1 Intuitive physics
Figure 5: Example images from [86] which are either challenging or easy for Humans, GPT4-V, or both in the binary tower stability judgement task for physical intuition.
A.2 Causal reasoning
Figure 6: Example images from [87] which are either challenging or easy for Humans, GPT4-V, or both in the number of blocks that will fall task for causal reasoning. Easy here refers to images where judgements have a low average distance to the that the ground truth. Hard refers to images where judgements have a high average distance to the ground truth.
A.3 Intuitive psychology
Figure 7: Example images from experiment 1B where the ratings given by GPT4-V match or diverge from human psychological intuition. A: GPT4-V answers 7, 5, 2, 7 for left cost, right cost, left reward, right reward. These values make sense given the path of the agent. It crosses straight through the left terrain, indicating that it might have a high cost associated with it and that the agent does not find the left care package rewarding. It then crosses through the yellow terrain to pick up the white care package, indicating that this package has a high reward associated with it. B: GPT4-V answers 5, 4, 2, 7 for left cost, right cost, left reward, right reward. These values are counterintuitive. The agent is seen crossing into the left terrain and picking up the left reward, which indicates that the left care package has a significant reward associated with it and that the left terrain should not incurr a large cost upon the agent. GPT4-V however assigns the left terrain a higher cost than the right terrain and also the left care package a lower reward than the right care package.
Figure 8: Example images from experiments 1A and 1B. A: GPT4-V answers 7, 7, 0 for inside cost, outside cost, top reward. That the care package has no value for the agent is sensible given the image, however we would expect the costs for inside and outside terrains to diverge, as the agent seemingly takes a detour to avoid crossing the inside terrain. B: GPT4-V answers 5, 5, 7 for bottom cost, top cost, top reward. The reward here should have a high value seeing as the agent takes a detour to collect the care package. However, the bottom terrain should be associated with a lower cost compared to the top terrain, as the agent is seen taking a detour to cross the bottom terrain instead of the top terrain after picking up the care package.
Prompts
Intuitive physics
For the intuitive physics experiment, we used a task from Lerer et al. [86]. For each trial, we asked the models three questions:
Question 1
What is the background color?
Question 2
How many blocks are in the image?
Question 3
Will this block tower fall? Give a boolean answer.
Causal reasoning
For the causality experiment, we used a task from Zhou et al. [87, 89]. For each trial, we again asked the models three questions:
Question 1
How many blocks are there in the image? You are only allowed to answer with a single number. No words allowed!
Question 2
How many of the red bricks would fall off the table if the dark grey brick wasn’t there? You are only allowed to answer with a single number between 0 and {num_blocks} corresponding to how many blocks would fall. No words allowed!
Question 3
How responsible is the dark grey brick for the red bricks staying on the table? You are only allowed to answer with a number on a scale from 0% (not at all responsible) to 100% (fully responsible). No words allowed!
Naive Utility Calculus
For the intuitive psychology experiment, we used three tasks from Jara-Ettinger et al. [88]. We again asked the model two descriptive questions in order to assess their basic comprehension of the scene:
Question 1
Please answer the following question with a single color name only: What color is the background of the central part of the image? You are only allowed to answer with a single color name!
Question 2
Please answer the following question with a number only: How many orange or white containers are in the image? You are only allowed to answer with a number!
For experiment 1A we first gave the models the following basic prompt, which was combined with different trial specific questions:
[…]
This task is about astronauts. The astronauts are exploring planets with alien terrains depicted with different colours and textures. Each astronaut has different skills, making each terrain more or less exhausting or easy for them to cross. All astronauts can ultimately cross all terrains, even if it’s exhausting. The astronauts land far from the base and have to walk there. In each image, the black circle on the left indicates where the astronaut landed. The base is on the middle right part of the image. Sometimes care packages are dropped from above and the astronauts can pick them up. There are two kinds of care packages depicted with an orange cylinder and a white cube. Each astronaut has different preferences and likes each kind of care package in different amounts. The astronauts don’t actually need the care packages. They can go straight to the base, or they can pick one up. You will see images of different astronauts with different skills and preferences travelling from their landing location to the home base. The astronauts always have a map. So they know all about the terrains and the care packages. Please answer the following question with a number only:
For experiment 1B we first gave the models the following basic prompt, which was again combined with different trial-specific questions:
[…]
This task is about astronauts. The astronauts are exploring planets with alien terrains depicted with different colours and textures. Each astronaut has different skills, making each terrain more or less exhausting or easy for them to cross. All astronauts can ultimately cross all terrains, even if it’s exhausting. Sometimes, the astronauts land far from the base and have to walk there. In each image, the black circle indicates where the astronaut landed. The base is in the center of the image. Sometimes care packages are dropped from above and the astronauts can pick them up. There are two kinds of care packages depicted with an orange cylinder and a white cube. Sometimes both care packages are identical. The astronauts cannot pick both care packages. Each astronaut has different preferences and likes each kind of care package in different amounts. The astronauts don’t actually need the care packages. They can go straight to the base, or they can pick one up. You will see images of different astronauts with different skills and preferences travelling from their landing location to the home base. The astronauts always have a map. So they know all about the terrains and the care packages. Please answer the following question with a number only:
Finally, for experiment 1C we first gave the models the following basic prompt:
[…]
This task is about astronauts. The astronauts are exploring planets with alien terrains depicted with different colours and textures. Each astronaut has different skills, making each terrain more or less exhausting or easy for them to cross. All astronauts can ultimately cross all terrains, even if it’s exhausting. The astronauts land far from the base and have to walk there. In each image, the black circle on the left indicates where the astronaut landed. The base is on the right part of the image. The path astronauts take from where they land to their base is indicated by a thick black line between the black circle on the left and the astronaut on the right. Sometimes care packages depicted by a blue cube on a black background are dropped from above and the astronauts can pick them up. Each astronaut has different preferences and likes each care package in different amounts. The astronauts don’t actually need the care packages. They can go straight to the base, or they can pick one up. You will see images of different astronauts with different skills and preferences travelling from their landing location to the home base. Your task is to judge how easy/exhausting it is for the astronaut in each image to cross each terrain, and how much they like each care package. The astronauts always have a map. So they know all about the terrains and the care packages. Please answer the following question with a number only:
Experiment 1A
Images Questions
0, 1
  1. 1.
    How easy is it for the astronaut in this image to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How much does the astronaut in this image like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
2, 3, 4, 5
  1. 1.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How much does the astronaut like the white care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!"
6, 7, 8, 9, 10
  1. 1.
    How easy is it for the astronaut to cross the purple terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
11, 12, 13, 14, 15
  1. 1.
    How easy is it for the astronaut to cross the purple terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the white care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
  4. 4.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!"
Experiment 1B
Images Questions
1, 2, 3, 4, 5, 6, 7
  1. 1.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How easy is it for the astronaut to cross the yellow terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
  4. 4.
    How much does the astronaut like the white care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
8, 9, 10
  1. 1.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How easy is it for the astronaut to cross the yellow terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
11, 12, 13, 14, 15, 16, 17
  1. 1.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How much does the astronaut like the orange care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the white care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!
Experiment 1C
Images Questions
All images
  1. 1.
    How easy is it for the astronaut to cross the yellow terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  2. 2.
    How easy is it for the astronaut to cross the pink terrain on a scale from 0 (extremely easy) to 10 (extremely exhausting)? You are only allowed to answer with a number!
  3. 3.
    How much does the astronaut like the blue care package on a scale from 0 (not at all) to 10 (a lot)? You are only allowed to answer with a number!

### 24 Jan 2024 v2

@@
-Have we built machines that think like people?
-Abstract
-A chief goal of artificial intelligence is to build machines that think like people.
-Yet it has been argued that deep neural network architectures fail to accomplish this.
-Researchers have asserted these models’ limitations in the domains of causal reasoning, intuitive physics, and intuitive psychology.
-Yet recent advancements, namely the rise of large language models, particularly those designed for visual processing, have rekindled interest in the potential to emulate human-like cognitive abilities.
-This paper evaluates the current state of vision-based large language models in the domains of intuitive physics, causal reasoning, and intuitive psychology.
-Through a series of controlled experiments, we investigate the extent to which these modern models grasp complex physical interactions, causal relationships, and intuitive understanding of others’ preferences.
-Our findings reveal that, while these models demonstrate a notable proficiency in processing and interpreting visual data, they still fall short of human capabilities in these areas.
-The models exhibit a rudimentary understanding of physical laws and causal relationships, but their performance is hindered by a lack of deeper insights—a key aspect of human cognition.
-Furthermore, in tasks requiring an intuitive theory of mind, the models fail altogether.
-Our results emphasize the need for integrating more robust mechanisms for understanding causality, physical dynamics, and social cognition into modern-day, vision-based language models, and point out the importance of cognitively-inspired benchmarks.
-Introduction
-Figure 1: Overview of domains tasks, approach, and models.
-A: Example images for the different experiments.
-Each experiment was taken from one of three cognitive domains: intuitive physics, causal reasoning, and intuitive psychology.
-B: General approach.
-For every query, an image was submitted to the model, and different questions were asked about the image, i.e.
-we performed visual question answering.C: Used multi-modal large language models and their size.
-People are quick to anthropomorphize, attributing human characteristics to non-human agents [1].
-This tendency, far from being a modern quirk, is deeply rooted in our collective psyche, as evidenced through our literary traditions.
-Consider E.T.A.
-Hoffmann’s “Der Sandmann” [2], where Nathaniel’s infatuation with Olimpia, a mere mechanical automaton with a limited repertoire of expressions – her most notable being a plaintive ’Ach, Ach!’ – serves as a poignant example.
-With further technological advances, this phenomenon, once confined to the realm of literature, has now manifested in reality.
-Already simple chatbots like ELIZA [3], which used simple pattern matching and substitution methodologies, gave users the illusion of human-like understanding and engagement [4].
-The tendency to anthropomorphize has only intensified with the advent of Large Language Models (LLMs) [5].
-These advanced AI systems represent a big leap from ELIZA both in complexity and capability.
-Unlike ELIZA’s simplistic pattern matching, LLMs apply deep learning techniques to generate text [6], learning from vast datasets to produce responses that can be startlingly human-like [7].
-Astonishingly, these models cannot only generate text.
-When scaled up to bigger training data and architectures, other, so-called “emergent abilities” appear [8, 9].
-The current models can, for example, pass the bar exam [10], write poems [11], compose music [12], and assist in programming and data analysis tasks [13].
-As a result, the line between human and machine capabilities is increasingly blurred [14, 15].
-People not only interact with these systems as if they were humans [16], but they also start to rely on them for complex decision-making [17], artistic creation [18], and personal interactions [19].
-It is, therefore, natural to ask: Have we built machines that think like people?
-Or are we, just like Nathaniel, projecting human qualities into rudimentary entities that are fundamentally different from us?
-Judging whether or not artificial agents can mimic human thought is at the core of cognitive science [20, 21].
-Therein, researchers have long debated the capabilities of artificially intelligent agents [22, 23, 24].
-In a seminal paper, Lake et al.
-[25] proposed core domains to consider when making such judgments.
-Published during the height of the deep learning revolution [26], the authors focused on domains that were easy for people but difficult for deep learning models: intuitive physics, causal reasoning, and intuitive psychology.
-Research on intuitive physics has studied how people perceive and interpret physical phenomena [27, 28, 29].
-Past work on this topic has emphasized that humans possess an innate ability to predict and understand the physical properties of objects and their interactions [30], even from a young age [31], a notion sometimes summarized as a “physics engine” in people’s heads [32].
-This understanding includes concepts such as gravity [33], inertia [34], and momentum [35].
-Some of the most canonical tasks in this domain involve testing people’s judgments about the stability of block towers [36, 37].
-These tasks have made their way into machine learning benchmarks [38, 39], where they are used to test the intuitive physical understanding of neural networks (see [40] for an overview of previous work on building models with human-like physical knowledge).
-Research on causal reasoning has studied how individuals infer and think about cause-effect relationships [41, 42, 43].
-Past work on this topic has proposed that humans possess an intuitive capacity to infer, understand, and predict causal relationships in their environment [44, 45, 46, 47], oftentimes described using Bayesian models of causal learning [48, 49].
-This cognitive ability encompasses recognizing patterns [50, 51], inferring causes from interventions [52, 53], and predicting future events based on hypothetical events [54].
-Canonical tasks in this domain often involve assessing individuals’ ability to infer causal relationships, for example, when judging the responsibility of one object causing other objects’ movement [55, 56].
-Causal reasoning remains a challenge, even for current machine learning approaches [57, 58].
-Research on intuitive psychology has explored how individuals infer, understand, and interpret social phenomena and mental states of other agents [59, 60].
-Past work on this topic has emphasized the concept that humans possess an inherent ability to infer and reason about the mental states [61, 62], intentions, and emotions of others, often referred to as a “theory of mind” [63, 64].
-This ability has been modeled as a Bayesian inference problem [65, 66, 67].
-Canonical tasks in this domain often involve assessing individuals’ capacity to predict actions based on understanding others’ perspectives or intentions, such as determining agents’ utility functions based on their actions in a given environment [68, 69].
-It is the subject of ongoing debates if modern algorithms show any form of intuitive psychology [70, 71, 72].
-Lake et al.
-argued that some of these abilities act as “start-up software,” because they constitute cognitive capabilities present early in development.
-Moreover, they proposed that these so-called “intuitive theories” [73, 74] need to be expressed explicitly using the calculus of Bayesian inference [75], as opposed to being learned from scratch, for example, via gradient descent.
-However, with the increase in abilities of current neural networks, in particular LLMs, we pondered: Can LLMs, in particular vision LLMs, sufficiently solve problems from these core domains?
-To address this question, we took canonical tasks from the domains of intuitive physics, causal reasoning, and intuitive psychology that could be studied by providing images and language-based questions.
-We submitted them to some of the currently most advanced LLMs.
-Our results showed that these models can, at least partially, solve these tasks.
-In particular, the largest currently available model, OpenAI’s Generative Pre-trained Transformer (GPT-4) managed to perform robustly above chance in two of the three domains.
-Yet crucial differences emerged.
-First, none of the models matched human-level performance in any of the domains.
-Secondly, none of the models fully captured human behavior, leaving room for domain-specific models of cognition such as the Bayesian models originally proposed for the tasks.
-Results
-We tested four different models on three core components for human-like intelligence as outlined by Lake et al.
-([25]; see Fig.
-1A).
-The models we used are vision large language models, which are multimodal models that integrate image processing capabilities into large language models ([76, 77]; see Fig.
-1B).
-These models allow users to perform visual question answering [78, 79]: users can upload an image and ask questions about it, which the model interprets and responds to accordingly.
-The first model was GPT-4 with Vision (GPT-4V), developed by OpenAI [80].
-This multimodal model extends the abilities of GPT-4 to analyze and interpret images, although the details of how the model accomplished this have not been made public.
-The second model is ADEPT’s Fuyu-8B [81], which is a decoder-only multimodal model based on the classic transformer architecture.
-Fuyu-8B stands out due to its simpler architecture and training procedure compared to other multimodal models.
-The third model is LLaMA-Adapter V2 [82], or short Adapter, which offers an alternative approach to enhance the capabilities of vision-language models by increasing learnable parameters, introducing an early fusion strategy for better visual integration, and employing a joint training approach for image-text and instruction-following data.
-The final model is Otter [83], which is based on OpenFlamingo [84], an open-sourced version of DeepMind’s Flamingo [85], and specifically designed for multimodal in-context instruction following.
-For testing the three core components, we used three tasks taken from the cognitive science literature that could be studied in vision LLMs via visual question answering.
-For intuitive physics, we asked models to judge the stability of different block towers, using stimuli originally proposed by Lerer and colleagues [86] and based on previous research by Battaglia et al.
-[36].
-For causal reasoning, we again used block towers and asked how many blocks would fall if certain blocks were removed or to judge the responsibility of certain blocks for the stability of the tower, adopting a design that Zhou and colleagues had previously used in human subjects [87].
-Finally, for intuitive psychology, we used a task where the models saw a picture of an agent’s path on a grid and then had to make inferences about the costs and rewards associated with the environment, taken from Jara-Ettinger and colleagues [88].
-For every task, we queried the visual reasoning abilities of the LLMs with tasks of increasing complexity (see Fig.
-1C).
-First, we asked about simple features of the shown images such as the background color or the number of objects shown.
-Afterward, we submitted questions taken from the cognitive science experiments.
-We report results based on comparisons to the ground truth as well as the different models’ matches to human data.
-Figure 2: Results for four different large language models for tasks of increasing complexity given images of real block towers from [86].
-We first ask for the background color in the image (A), then the number of blocks in the image (B), and finally a binary stability rating for the block towers (C).
-The last plot shows the match to human stability ratings as determined by the Pearson correlation (D).
-Errors bars in plots A - C are given by the standard deviation of a binomial distribution.
-The error bars in plot D are given by the 95% confidence interval for the correlation coefficient.
-Intuitive physics
-To test the intuitive physics capabilities of the different LLMs, we used photographs depicting wooden block towers from Lerer et al.
-[86] (see Fig.
-5 in the Appendix for an example).
-These images mirror stimuli that developmental psychologists use to study the development of intuitive physics in infants.
-We used these images to test the models in increasingly complex tasks, starting with determining the background color of a given image (1), counting the number of colored blocks in the image (2), and giving a binary stability judgment of the depicted block tower (3).
-All four models were able to correctly perform the first and easiest task: they all achieved almost perfect accuracy in determining the background color of the images (see Fig.
-2A).
-In the second task, the performance of most models deteriorated, with only GPT-4V correctly determining the number of blocks in all images (see Fig.
-2B).
-It is important to note that the first two tasks are fairly trivial for humans and we would expect human performance to be at 100% (the background color is always white and images featured either 2, 3, or 4 blocks).
+bibliography
 […]
-[…]
-Causal reasoning
-To test the causal reasoning capabilities, we used synthetic images again depicting block towers from Zhou et al.
-[89, 87] (see Fig.
-6 in the Appendix for an example).
-Here, the images showed static scenes of block towers that were stable but might collapse if one of the blocks was removed.
-Figure 3: Results for analyses using images from causal reasoning experiments taken from [89].
-We first ask for the number of blocks in the image, then the number of blocks that would fall if a specific block is removed, and finally a rating between 0 and 100 for how responsible a specific block is for the stability of the tower.
-For the responsibility ratings, all LLMs except for GPT-4V give constant ratings: Fuyu always responds with 100, while Otter and LLaMA-Adapter V2 always respond with 50.
-Error bars in plots A - C are given by the standard error of the mean, while the error bar plot D is given by the 95% confidence interval of the correlation coefficient.
-Again, we started by asking the models to count the blocks in the image (1), we continued by querying the models for the number of blocks that would fall if a specific block was removed from the scene (2), and finally, we asked the models to rate the responsibility of a specific block for the stability of the other blocks (3).
-For the second task, we established a baseline performance represented by a horizontal line in Fig.
-3A-C, which corresponds to a random agent.
-This random agent used a simple strategy: it gave the mean between 0 and the number of blocks in each image as its’ prediction, essentially behaving like a uniform distribution over the possible number of blocks that could fall.
-The images in this task displayed a larger number of blocks (ranging from 6 to 19), which made the basic counting task significantly more challenging than in the previous section.
-Models’ responses approximated the ground truth, albeit rarely matching it exactly.
-Therefore, we report the mean absolute distance to the ground truth instead of the percentage of correct answers (see Fig.
-3A).
-The models’ performance highlighted the challenging nature of this task, with the best performing model (GPT4-V) still being on average  2 blocks off.
-[…]
-[…]
-Intuitive psychology
-To test the intuitive psychology of the different LLMs we used synthetic images depicting an astronaut on a colored background from Jara-Ettinger et al.
-[88] (see Fig.
-5 in the Appendix for an example).
-The experiment consisted of three parts which differed slightly in their layout and components.
-Depending on the experiment, the images featured either one or two different terrains (indicated by different background colors and textures) as well as either one or two different care packages.
-The astronaut was shown with a path that led from a starting point to a base; the astronaut could collect care packages along the way.
-Depending on which terrain the astronaut crossed or which care package they chose to pick up or not, it was possible to infer the costs associated with the different terrains and the rewards associated with the different care packages.
-Figure 4: Results on tasks for intuitive psychology taken from [88].
-Again, we first ask for the background color (A).
-Models are then asked to make inferences about the costs and rewards in an environment depending on the path an agent has taken (B and C).
-Correlation values for Fuyu and LLaMA-Adapter V2 are missing as they always responded with constant ratings for either cost or reward questions.
-Error bars in plot A are given by the standard deviation of a binomial distribution, while the error bars in plots B and C are given by the 95% confidence interval for the correlation coefficient.
-[…]
-Discussion
-We started by asking whether, with the rise of modern large language models, researchers have created machines that – at least to some degree – think like people.
-To address this question, we took four recent multi-modal large language models and probed their abilities in three core cognitive domains: intuitive physics, causal reasoning, and intuitive psychology.
-In intuitive physics, the models managed to solve some of the given tasks and showed a medium match with human data.
-Similarly, in a causal reasoning task, some models, in particular, GPT-4V, performed well and showed a medium match with human data.
-Finally, in an intuitive psychology task, none of the models performed well and none of them showed a reasonable match with human data.
-Thus, an appropriate answer to the question motivating our work would be “No.”, or – perhaps more optimistically – “Not quite.”
-Limitations and Future Work
-Although we have tried our best to give all models a fair chance and set up the experiments in a clean and replicable fashion, some shortcomings remain that should be addressed in future work.
-First of all, we have only tested a handful of multi-modal models on just three cognitive domains.
-While we believe that the used models and tasks provide good insights into the state-of-the-science of LLMs’ cognitive abilities, future studies should look at more domains and different models to further tease apart when and why LLMs can mimic human reasoning.
-For example, it would be interesting to see if scale is the only important feature influencing model performance [90, 91].
-Currently, our evidence suggests that even smaller models, for example Fuyu with its 8 billion parameters, can sometimes perform as well as GPT-4V in some tasks.
-Another shortcoming of the current work is the simplicity of the used stimuli.
-While the block towers used in our first study were deliberately designed to be more realistic [86] than commonly used psychological stimuli [36], this was not true for the experiments in the other two domains.
-For the intuitive psychology experiments, in particular, we would expect the models to perform better if the stimuli contained more realistic images of people, which has been shown to work better in previous studies [92].
-Interestingly, using more realistic stimuli can also change people’s causal judgments [93]; how realistic stimuli used in cognitive experiments should be, remains an open question [94].
-On a related point, we only used static images in our current experiments, which severely limits the breadth and level of detail of the questions we could ask.
-For example, some of the most canonical tasks investigating people’s causal reasoning abilities involve videos of colliding billiard balls [55].
-As future large language models will likely be able to answer questions about videos [95], these tasks represent the next frontier of cognitively-inspired benchmarks.
-For the comparisons to human data, we currently used the participant data collected in the original studies and assessed the correspondence between models and this data via correlation coefficients.
-Future work could expand on this approach by collecting new data from human subjects choosing which of the model’s judgments they prefer.
-This could lead to a more detailed comparison, similar to what has been proposed to discriminate among deep learning models for human vision [96] and language [97].
-A crucial weakness of most studies using large language models is that they can be sensitive to specific prompts [98, 99, 100].
-While we have attempted to use prompts that elicited good behavior, thereby giving LLMs a chance to perform well, future work could try to further optimize these prompts using available methods [101, 102, 103], while also assessing how the models respond to paraphrased versions of the same tasks.
-While it could be possible to further engineer the used prompts, as well as try out several other ways of phrasing the same prompt, we believe that our current approach was sufficient to showcase these models’ abilities.
-Finally, we applied all models out of the box and without additional fine-tuning.
-Future studies could attempt to fine-tune multi-modal LLMs to better align with cognitive data [104] and assess if this improves their reasoning abilities more generally.
-Related work
-We are not the first to assess LLMs’ reasoning abilities [105, 106, 107].
-Previous studies have focused, among others, on testing LLMs’ cognitive abilities in model-based planning [108], analogical reasoning tests [109], exploration tasks [110], systematic reasoning tests [111, 112], psycholinguistic completion studies [113], and affordance understanding problems [114].
-In this sense, our contribution can be seen as a part of a larger movement in which researchers use methods from the behavioral sciences to understand black box machine learning models [115, 116, 117].
-However, most of the previous studies did not investigate multi-modal LLMs but rather remained in the pure language domain.
-Although there are recent attempts to investigate vision LLMs cognitive features, including their reaction to visual illusions [118] as well as how they solve simple intelligence tasks [119], we are the first to investigate the proposed core components of cognition in these models.
-Previous work has also looked at how LLMs solve cognitive tasks taken from the same domains that we have looked at.
-In intuitive physics, Zečević et al.
-[120] found that LLMs performed poorly in a task using language descriptions of physical scenarios.
-Zhang and colleagues [121] extracted programs from text produced by large language models to improve their physical reasoning abilities.
-Finally, Jassim and colleagues [122] proposed a novel benchmark for evaluating multimodal LLMs’ understanding of situated physics.
-In causal reasoning, Binz and Schulz [123] showed that GPT-3 failed at simple causal reasoning experiments, while Kosoy et al.
-[124] showed that LLMs cannot learn human-like causal over-hypotheses.
-In research on intuitive psychology, Kosinsky argued that theory of mind might have emerged in LLMs [71] which has been criticized other researchers [72].
-Akata et al.
-showed that GTP-4 plays repeated games very selfishly and could not pick up on simple conventions such as alternating between options [19].
-Finally, Gandhi and colleagues [125] proposed a framework for procedurally generating Theory of Mind evaluations and found that GPT4’s abilities mirror human inference patterns, though less reliable, while all other LLMs struggled.
-Many of the past studies on LLMs have fallen risk of appearing in new models’ training set.
-Recent work has recognized this issue and, in turn, evaluated language models on many problem variations to minimize training set effects [126].
-Our work differs from these approaches as current models could not have just memorized solutions to the given problems because these problems require deep reasoning and are rarely ever published with the ground truth attached.
-Conclusion
-A major plot twist in E.T.A Hoffmann’s “Der Sandmann” is that Nathaniel might have been a machine himself, which explains why he fell for Olimpia.
-This metaphor also relates to our anthropomorphization of LLMs: since LLMs are trained on human-generated data, their behaviors will always reflect our behaviors and biases.
-However, this reflection is sharpening, and modern neural network architectures have become more human-like.
-One of the other domains emphasized by Lake et al.
-was the ability to reason compositionally.
-Recent attempts have shown that neural networks can perform compositional reasoning if trained appropriately [127].
-Similarly, our current work has shown that multimodal LLMs have come a long way, showing some correspondence to human behavior and often performing above chance.
-Moreover, machine learning researchers have put forward various ideas about how to close the remaining gap between humans and machines [128], including self-supervised learning [129], translating from natural into probabilistic languages [130], or grounding LLMs in realistic environments [131].
-This continuous evolution in models’ capabilities necessitates a reevaluation of the metaphors and tools we use to understand them.
-We believe that cognitive science can offer tools, theories, and benchmarks to evaluate how close we have come to building machines that think like people.
-Methods
-Code
-The open-source models were installed per the instructions on their related Github or Huggingface repositories and evaluated on a Slurm-based cluster with a single A100.
-For the results reported as GPT-4V, we used the public ChatGPT interface and the OpenAI API, specifically the November 2023 release of gpt4-vision-preview model which is available via the completions endpoint.
-Code for replicating our results is available on GitHub (github.com/lsbuschoff/multimodal).
-All models were evaluated in Python using PyTorch [132].
-Additional analyses were carried out using NumPy [133], Pandas [134], and SciPy [135].
-Matplotlib [136] and Seaborn [137] were used for plotting.
-Models
-GPT4-V
-We initially queried GPT-4V through the ChatGPT interface, since the OpenAI API was not publicly available at the outset of this project.
-The Intuitive Psychology task responses were collected using the gpt4-vision-preview model variant after its November 2023 release in the API.
-We set the maximum number of generated tokens for a given prompt to 1 to get single numerical responses.
-All other parameters were set to their default values.
-Note that this model does not currently feature an option for manually setting the temperature, and the provided documentation does not specify what the default temperature is.
-FUYU
-Fuyu is an 8B parameter multi-modal text and image decoder-only transformer.
-We used the Huggingface implementation with standard settings and without further finetuning (available here).
-The maximum number of generated tokens was set to 8 and responses were parsed by hand.
-ADAPTER
-We here use the multi-modal version of LLaMA-Adapter V2, which adds adapters into LLaMA’s transformer in order to turn it into an instruction-following model.
-We used the GitHub implementation with standard settings and again without further finetuning (available here).
-The maximum number of generated tokens was left at 512 and responses were parsed by hand.
-OTTER
-Otter is a multi-modal LLM that supports in-context instruction tuning and it is based on the OpenFlamingo model.
-We used the Huggingface implementation (available here), again with standard settings and without finetuning.
-The maximum number of generated tokens was left at 512 and responses were parsed by hand.
-Datasets
-Intuitive physics
-We tested the intuitive physical understanding of the models using images from Lerer et al.
-[86].
-The photos depict a block tower consisting of colored wooden blocks in front of a white fabric (see Fig.
-5 for an example).
-The images are of size 224 x 244.
-In the data set, there are a total of 516 images of block towers.
-We tested the models on 100 randomly drawn images.
-We first tested the models on their high-level visual understanding of the scenes: we tasked them with determining the background color and the number of blocks in the image.
-In order to test their physical understanding, we tested them on the same task as the original study: we asked them to give a binary rating on the stability of the depicted block towers.
-For the first two tasks, we calculated the percentage of correct answers for each of the models.
-For the third task, we used the Pearson correlation to determine the degree of similarity between the models and human subjects.
-Causal reasoning
-For the causality part we used images from Zhou et al.
-[87].
-The images show artificial block stacks of red and gray blocks on a black table (see Fig.
-6 for an example).
-The data set consists of 42 images on which we tested all models.
-We again first tested the models on their high-level visual understanding of the scene and therefore tasked them with determining the number of blocks in the scene.
-The ground truth number of blocks in the scenes ranged from 6 to 19.
-Since this task is rather challenging due to the increased number of blocks, we do not report the percentage correct as for the intuitive physics data set but the mean over the absolute distance between model predictions and the ground truth for each image (see Fig.
-3A).
-To test the causal reasoning of the models we adopted the tasks performed in the original study [89].
-We asked models to infer how many red blocks would fall if the gray block was removed.
-We again report the absolute distance between model predictions and the ground truth for each image (see Fig.
-3B).
-We calculate a random baseline which uses the mean between 0 and the number of blocks for each specific image as the prediction.
-We also ask the models for a rating between 0 and 100 for how responsible the gray block is for the stability of the tower.
-For both, the number of blocks that would fall if the gray block was removed, and its’ responsibility for the stability of the tower, we calculate the mean Pearson correlation to human subjects from the original study (see Fig.
-3C).
-Intuitive psychology
-To test the intuitive psychology of the different LLMs, we used stimuli from Jara-Ettinger et al.
-[88].
-This part consisted of three different experiments each consisting of 16, 17, and 14 images showing a 2D depiction of an astronaut and care packages in different terrains (see Fig.
-5 for an example).
-In order to check their high-level understanding of the images, we again asked the models to determine the background color of the images.
-Since this background color is not uniform, we counted both “Pink” and “Purple” as correct answers.
-We report the percentage of correct answers for the background color in Figure 2A.
-In accordance with the original study, analyses for the intuitive psychological capabilities of the models are split into cost questions (passing through a terrain is associated with a cost for the agent) and reward questions (collecting a care package yields some sort of reward for the agent).
-We pooled cost and reward questions over all three experiments and reported the mean Pearson correlation with both humans and a heuristic outlined in [88] (see Figs.
-2B and 2C).
-This heuristic calculates the costs and rewards associated with the environment from the amount of time an agent spends in each terrain and which care package it collects.
-Acknowledgements
-This work was supported by the Max Planck Society, the Volkswagen Foundation, the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A, and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2064/1 – 390727645.
-Author contributions statement
-All authors conceived the experiments. L.B.S. and E.A. conducted the experiments. All authors analysed the results. All authors wrote the manuscript.
-Appendix A Example trials
-A.1 Intuitive physics
-Figure 5: Example images from [86] which are either challenging or easy for Humans, GPT4-V, or both in the binary tower stability judgement task for physical intuition.
-A.2 Causal reasoning
-Figure 6: Example images from [87] which are either challenging or easy for Humans, GPT4-V, or both in the number of blocks that will fall task for causal reasoning.
-Easy here refers to images where judgements have a low average distance to the that the ground truth.
-Hard refers to images where judgements have a high average distance to the ground truth.
-A.3 Intuitive psychology
-Figure 7: Example images from experiment 1B where the ratings given by GPT4-V match or diverge from human psychological intuition.
-A: GPT4-V answers 7, 5, 2, 7 for left cost, right cost, left reward, right reward.
-These values make sense given the path of the agent.
-It crosses straight through the left terrain, indicating that it might have a high cost associated with it and that the agent does not find the left care package rewarding.
-It then crosses through the yellow terrain to pick up the white care package, indicating that this package has a high reward associated with it.
-B: GPT4-V answers 5, 4, 2, 7 for left cost, right cost, left reward, right reward.
-These values are counterintuitive.
-The agent is seen crossing into the left terrain and picking up the left reward, which indicates that the left care package has a significant reward associated with it and that the left terrain should not incurr a large cost upon the agent.
-GPT4-V however assigns the left terrain a higher cost than the right terrain and also the left care package a lower reward than the right care package.
-Figure 8: Example images from experiments 1A and 1B.
-A: GPT4-V answers 7, 7, 0 for inside cost, outside cost, top reward.
-That the care package has no value for the agent is sensible given the image, however we would expect the costs for inside and outside terrains to diverge, as the agent seemingly takes a detour to avoid crossing the inside terrain.
-B: GPT4-V answers 5, 5, 7 for bottom cost, top cost, top reward.
-The reward here should have a high value seeing as the agent takes a detour to collect the care package.
-However, the bottom terrain should be associated with a lower cost compared to the top terrain, as the agent is seen taking a detour to cross the bottom terrain instead of the top terrain after picking up the care package.
+Supplementary information
 Prompts
 Intuitive physics
-For the intuitive physics experiment, we used a task from Lerer et al. [86]. For each trial, we asked the models three questions:
+For the intuitive physics experiment, we used a task from Lerer et al. [lerer2016learning]. For each trial, we asked the models three questions:
 Question 1
@@
 Causal reasoning
-For the causality experiment, we used a task from Zhou et al. [87, 89]. For each trial, we again asked the models three questions:
+For the causality experiment, we used a task from Zhou et al. [zhou2023mental, gerstenberg2017faulty]. For each trial, we again asked the models three questions:
 Question 1
@@
 For the intuitive psychology experiment, we used three tasks from Jara-Ettinger et al.
-[88].
+[jara2020naive].
 We again asked the model two descriptive questions in order to assess their basic comprehension of the scene: